OpenStudy (anonymous):

How many millimeters of mercury is 2.64 atmospheres equivalent to? 2.01 x 10^3 mm Hg 3.47 x 10^−3 mm Hg 2.88 x 10^2 mm Hg 2.64 mm Hg

OpenStudy (anonymous):

Solve for x: \(\sf 1 ~atm=760~mm~Hg\) \(\sf 2.64~atm=x~~mm~Hg\)

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

Cross multiply them. \(\sf 1\times x =760 \times 2.64 \Longrightarrow x \approx2.01\times 10^3~mm~Hg\)